# MedTutor: A Retrieval-Augmented LLM System for Case-Based Medical Education

### v1

@@ 1 Introduction
The training of medical residents is an intensive learning process, built upon the foundation of studying and interpreting thousands of case reports. Residents routinely engage with clinical cases through discussions with peers and mentors, analyzing findings and differential diagnoses to deepen their understanding. However, while direct feedback from attending physicians is invaluable, the process of finding relevant educational material and supporting evidence for specific cases is often time-consuming and inconsistent (Rogers et al., 2019; Daniel et al., 2020). The sheer volume of medical literature and the challenge of identifying pertinent resources for each case can limit the depth of learning that residents achieve from their clinical experiences (Anderson and Anderson, 2019; Bednarczyk et al., 2014). There exists a significant opportunity to augment this traditional learning process with AI tools that can efficiently retrieve and synthesize educational content from clinical cases, drawing upon vast archives of medical knowledge. LLMs present a promising avenue for this augmentation, but their application in high-stakes medical domains is fraught with challenges, most notably the risk of factual inaccuracy (or hallucination) and the use of outdated knowledge (Abd-alrazaq et al., 2023; Li et al., 2023; Xie et al., 2023).
@@ 2.1 LLMs and RAG in the Medical Domain
The application of LLMs in medicine has shown immense promise. General-purpose foundation models have demonstrated impressive capabilities on standardized medical exams and complex diagnostic problems (Nori et al., 2023; Singhal et al., 2023). This has fueled a broader vision for generalist biomedical AI that can assist with a wide range of clinical tasks (Tu et al., 2023). However, the “black-box” nature of these models and their potential for factual errors or "hallucinations" remain significant barriers to clinical adoption, necessitating robust evaluation frameworks (Huang et al., 2024; Li et al., 2023; Xie et al., 2023).
To mitigate these risks, RAG has emerged as a key paradigm for building dependable clinical tools (Lewis et al., 2021). By grounding LLM outputs in external, verifiable evidence from reliable medical literature, RAG provides a pathway to trustworthy AI. A recent perspective in Nature Medicine strongly advocates for RAG as a prerequisite for the responsible deployment of generative AI in healthcare (Yang et al., 2024). The field is now maturing to a point where standardized benchmarks for medical RAG are being established, allowing for more rigorous evaluation of these systems (Xiong et al., 2024b). Our work contributes to this growing body of literature by presenting a novel RAG-based system specifically designed for medical education, a domain where accuracy and reliability are paramount.
@@ 2.2 LLMs for Medical Education
LLMs show promise in generating medical exam content, though concerns about accuracy necessitate expert oversight Zhu et al. (2024). Integrating RAG improves reliability by grounding output in external sources, with studies reporting notable gains in question-answering accuracy using medical textbooks Chen et al. (2025); Wang et al. (2023). Benchmarks like MIRAGE further validate RAG’s role in medical QA tasks Xiong et al. (2024a).
For resident training, LLMs can assess skills and provide feedback, but expert review remains vital Atsukawa et al. (2024). Systems enabling citation generation enhance factuality Wang et al. (2025), while evaluation frameworks like LLM-as-a-Judge offer scalability despite only moderate alignment with human judgment Zheng et al. (2025). New approaches continue to embed evidence-based medicine principles into RAG pipelines for clinically accurate educational content Lu et al. (2025).
@@ 3.1 The MedTutor Pipeline Stages
Hybrid Evidence Retrieval: For each search query, the system performs a hybrid retrieval process in parallel described below: (1) Local retrieval for textbook material: Textbooks and notes are essential resources for medical education. In our MedTutor system, we first apply OCR to a radiology textbook (Dahnert and Ovid Technologies, 2017) using the mistral-ocr-2503 model, then segment and index the material by page. We generate dense embeddings for these materials with the Qwen3-Embedding-8B model, which has demonstrated state-of-the-art performance in embedding and retrieval tasks on the MTEB benchmark (Muennighoff et al., 2022) among models of comparable size. These embeddings are stored in a pre-computed vector database for subsequent queries. For local database search, we employ a bi-encoder architecture to generate dense vector representations for both the query and the pre-indexed textbook pages, subsequently identifying the most relevant page using cosine similarity. (2) Retrieval using academic APIs: Some case reports are more specialized or rare, requiring retrieving knowledge from latest academic literature. Therefore, we also employ retrieval from academic search engines. We use PubMed and Semantic Scholar APIs, two commonly used and freely available scholarly systems, to fetch the latest relevant research papers. To prevent rate-limiting, API calls are managed by an asyncio.Semaphore. If pre-fetched results for the queries are available, this step is skipped to improve efficiency.

### v2

@@ 1 Introduction
The training of medical residents is an intensive learning process, built upon the foundation of studying and interpreting thousands of case reports. Residents routinely engage with clinical cases through discussions with peers and mentors, analyzing findings and differential diagnoses to deepen their understanding. However, while direct feedback from attending physicians is invaluable, the process of finding relevant educational material and supporting evidence for specific cases is often time-consuming and inconsistent (Rogers et al., 2019; Daniel et al., 2020). The sheer volume of medical literature and the challenge of identifying pertinent resources for each case can limit the depth of learning that residents achieve from their clinical experiences (Anderson and Anderson, 2019; Bednarczyk et al., 2014). There exists a significant opportunity to augment this traditional learning process with AI tools that can efficiently retrieve and synthesize educational content from clinical cases, drawing upon vast archives of medical knowledge. LLMs present a promising avenue for this augmentation, but their application in high-stakes medical domains is fraught with challenges, most notably the risk of factual inaccuracy (or hallucination) and the use of outdated knowledge (Abd-alrazaq et al., 2023; Li et al., 2023; Xie and Wang, 2023).
@@ 2.1 LLMs and RAG in the Medical Domain
The application of LLMs in medicine has shown immense promise. General-purpose foundation models have demonstrated impressive capabilities on standardized medical exams and complex diagnostic problems (Nori et al., 2023; Singhal et al., 2023). This has fueled a broader vision for generalist biomedical AI that can assist with a wide range of clinical tasks (Tu et al., 2023). However, the “black-box” nature of these models and their potential for factual errors or "hallucinations" remain significant barriers to clinical adoption, necessitating robust evaluation frameworks (Huang et al., 2024; Li et al., 2023; Xie and Wang, 2023).
To mitigate these risks, RAG has emerged as a key paradigm for building dependable clinical tools (Lewis et al., 2021). By grounding LLM outputs in external, verifiable evidence from reliable medical literature, RAG provides a pathway to trustworthy AI. A recent perspective in Nature Medicine strongly advocates for RAG as a prerequisite for the responsible deployment of generative AI in healthcare (Yang et al., 2024). The field is now maturing to a point where standardized benchmarks for medical RAG are being established, allowing for more rigorous evaluation of these systems (Xiong et al., 2024). Our work contributes to this growing body of literature by presenting a novel RAG-based system specifically designed for medical education, a domain where accuracy and reliability are paramount.
@@ 2.2 LLMs for Medical Education
LLMs show promise in generating medical exam content, though concerns about accuracy necessitate expert oversight Zhu et al. (2025). Integrating RAG improves reliability by grounding output in external sources, with studies reporting notable gains in question-answering accuracy using medical textbooks Chen et al. (2025); Wang et al. (2025b). Benchmarks like MIRAGE further validate RAG’s role in medical QA tasks Xiong et al. (2024).
For resident training, LLMs can assess skills and provide feedback, but expert review remains vital Atsukawa et al. (2025). Systems enabling citation generation enhance factuality Wang et al. (2025a), while evaluation frameworks like LLM-as-a-Judge offer scalability despite only moderate alignment with human judgment Zheng et al. (2025). New approaches continue to embed evidence-based medicine principles into RAG pipelines for clinically accurate educational content Lu et al. (2025).
@@ 3.1 The MedTutor Pipeline Stages
Hybrid Evidence Retrieval: For each search query, the system performs a hybrid retrieval process in parallel described below: (1) Local retrieval for textbook material: Textbooks and notes are essential resources for medical education. In our MedTutor system, we first apply OCR to a radiology textbook (Dähnert, 2017) using the mistral-ocr-2503 model, then segment and index the material by page. We generate dense embeddings for these materials with the Qwen3-Embedding-8B model, which has demonstrated state-of-the-art performance in embedding and retrieval tasks on the MTEB benchmark (Muennighoff et al., 2023) among models of comparable size. These embeddings are stored in a pre-computed vector database for subsequent queries. For local database search, we employ a bi-encoder architecture to generate dense vector representations for both the query and the pre-indexed textbook pages, subsequently identifying the most relevant page using cosine similarity. (2) Retrieval using academic APIs: Some case reports are more specialized or rare, requiring retrieving knowledge from latest academic literature. Therefore, we also employ retrieval from academic search engines. We use PubMed and Semantic Scholar APIs, two commonly used and freely available scholarly systems, to fetch the latest relevant research papers. To prevent rate-limiting, API calls are managed by an asyncio.Semaphore. If pre-fetched results for the queries are available, this step is skipped to improve efficiency.